# A Survey on Medical Large Language Models: Technology, Application, Trustworthiness, and Future Directions

### v1

@@ III-C Evaluations for Med-LLMs
III-C1 Quantitative Evaluations
In this section, we summarize some commonly used metrics for NLP tasks. As shown in Table IV, different tasks are evaluated via various metrics, covering accuracy, F1 score, and perplexity. Here, we introduce the details of some important metrics.
Perplexity. Perplexity (PPL) is a standard that evaluates how well a probability model can predict a sample. When applied to language models like GPT, it represents the exponential average negative log-likelihood of a sequence. In essence, a lower perplexity score suggests that the model has a higher certainty in its predictions.
@@ III-C1 Quantitative Evaluations
BLEU. BLEU calculates the similarity between a candidate translation and one or more reference translations based on n-gram overlap. The idea is that a good translation should contain many n-grams (sequences of words) that are also present in the reference translations. Mathematically, the BLEU score is computed as follows:
@@
ROUGE. ROUGE scores compare the overlapping units (n-grams, skip-bigrams, or longest common subsequences) between the generated text and the reference. There are many ROUGE variants to measure the overlaps from different viewpoints, such as ROUGE-N, ROUGE-L, ROUGE-W, and ROUGE-SU. Here, we only give a detailed explanation for the most basic ROUGE-N. ROUGE-N is a metric that quantifies the overlapping N-grams between a candidate summary and a reference summary. Let $C$ be the set of all unigrams (or bigrams for ROUGE-2) in the candidate summary and $R$ be the set of all unigrams (or bigrams) in the reference summary. ROUGE-1 can be calculated as:
$P_{\text{ROUGE-1}}$ $=\frac{\sum_{w\in C}\text{Count}{R}(w)}{|C|},$ (3)
$R_{\text{ROUGE-1}}$ $=\frac{\sum_{w\in C}\text{Count}{R}(w)}{|R|},$
@@ III-C3 Evaluation Benchmark for LLMs’ Clinical Capabilities
MedBench [17]. MedBench is a comprehensive evaluation benchmark for the medical domain, which consists of 40,041 questions about various branches of medicine including authentic examinations and clinical reports. The main components of MedBench involve: Chinese Medical Licensing Examination, Resident Standardization Training Examination, Doctor In-Charge Qualification Examination, and real-world clinic cases encompassing examinations, diagnoses, and treatments. Extensive experiments provided the following findings: (1) Chinese medical LLMs (e.g., HuaTuoGPT [167]) underperform on the MedBench, indicating the importance of clinical knowledge and diagnostic precision. (2) Some general LLMs (e.g., GPT 4) surprisingly convey considerable clinical knowledge.
@@
AutoEval [29]. AutoEval is an automated evaluation framework to evaluate the realistic capabilities of LLMs, i.e., working as virtual doctors for multi-turn dialogue. The clinical demands for consultation tasks require LLMs to (1) consider the factors they do not know, (2) query about missing medical factors of patients, and (3) generate diagnosis and treatment results. To this end, AutoEval is to reformulate USMLE into medical multiple-choice questions. Then, several metrics are designed to indicate the ability of doctor LLMs in multi-turn consultation, i.e., Medical Information Coverage Rate, Accuracy of the Final Task, and Average Turn and Average Length.
@@
LLM-Mini-CEX [30]. Based on MiniCEX [181, 182], an evaluation criterion called LLM-specific Mini-CEX is established to effectively assess the clinical capabilities of LLMs. Then, a patient simulator is developed to simulate the conversations with LLMs, where ChatGPT is utilized to automatically evaluate diagnosis dialogues.
@@
MedGPTEval [31]. MedGPTEval contains Chinese medical datasets and public benchmarks. The evaluation metrics are designed based on a comprehensive literature review, including medical professional capabilities, social comprehensive capabilities, contextual capabilities, and computational robustness, with 16 detailed indicators. These metrics have been optimized by a Delphi method using 5 experts in medicine and engineering [31]. Then, 3 clinical experts manually construct a set of medical datasets for interactions, including 7 case reports and 27 medical dialogues.
@@
LLM-Human Evaluation [183]. LLMs have exhibited superior performance on unseen tasks, even only trained on the specific-task instructions. The study [183] investigates that if such an ability can be utilized as an alternative to human evaluation. Both human and LLM evaluations are conducted for two NLP tasks: adversarial attacks and open-ended story generation. The experiments indicate that LLM evaluation exhibits consistent results with human-level evaluation. Besides, LLM evaluation provides more stable results over different task instructions and the sampling algorithms for generation.
@@
RJUA-SPs [32]. The RJUA-SPs evaluation approach contains three basic elements: (1) Metric. Professional clinical practice pathways are exploited to define the clinical capabilities that a doctor should possess, named LLM-specific clinical pathway (LCP). (2) Data. Standardized Patients (SPs) are utilized to guide the data collection, which aims to well maintain the completeness of the evaluation. (3) Algorithm. A multi-agent framework is formulated to simulate the interactive environment between SPs and a doctor agent. A retrieval-augmented algorithm is specially designed to measure whether the behaviors of a LLM doctor are consistent with LCP.
III-D Specific Med-LLMs
@@
Existing LLMs in the medical field, such as Med-PALM, Codex-Med, and MedAlpaca, have contributed to advancements in healthcare with their unique design objectives, architectures, and functionalities. These models collectively drive the development of Med-LLMs and enhance their application in healthcare, providing robust support for future medical practice and research.
HuatuoGPT [167]. HuatuoGPT is a Chinese LLM for medical consultation purposes. Based on a reinforcement learning from AI feedback approach, a reward model is trained to align LLMs with both the distilled and real-world data. Three QA datasets are used to indicate the effectiveness of HuatuoGPT: cMedQA2, webMedQA, and Huatuo-26M. The evaluation metrics contain BLEU, GLEU, ROUGE, and DISTINCT. GPT-4 is utilized to indicate the quality of the generated outputs.
ClinicalT5 [33]. ClinicalT5 is a T5-based generative language model, which is specifically pre-trained on clinical text to address the linguistic challenges of clinical notes, posed by utilizations of professional terminology, abbreviations, and domain-specific writing patterns. ClinicalT5 aims to bridge the gap between general LLMs and the specialized nature of clinical text. The model is pre-trained using the span-mask denoising objective on a large corpus of clinical notes from the MIMIC-III database. Furthermore, ClinicalT5 is evaluated on real-world clinical applications, such as predicting patients’ 30-day readmission risk, 30-day mortality risk, and 1-year mortality risk.
ClinicalGPT [34]. ClinicalGPT is a specific LLM for clinical scenarios. By incorporating extensive and diverse real-world data in the training process, such as medical records, domain-specific knowledge, and multi-round dialogue consultations, ClinicalGPT is well-equipped to handle multiple clinical tasks. The training datasets for ClinicalGPT includes cMedQA2, cMedQA-KG, MD-EHR, MEDQA-MCMLE, and MedDialog, covering a wide range of medical information, such as question-answering pairs, electronic health records, and doctor-patient conversations.
ChiMed-GPT [6]. ChiMed-GPT is a medical LLM explicitly designed for clinical scenarios in Chinese language. The model is built upon Ziya-13B-v2, a general domain Chinese LLM, and undergoes a comprehensive training regime that includes PT, SFT, and RLHF. This holistic approach ensures that CHIMED-GPT not only captures domain-specific knowledge but also aligns with human preferences and adapts to various scenarios.
BioGPT [35]. BioGPT is a generative pre-trained language model specifically designed for biomedical text generation. Inspired by the success of GPT-like models, BioGPT aims to improve the weak generation ability for biomedical domain-specific language models, which primarily focus on BERT-like architectures. BioGPT adopts the GPT-2 model [184] as its backbone and it is pre-trained from scratch on PubMed abstracts. The model’s vocabulary is also constructed using byte pair encoding on the pre-training corpus, ensuring that it is well-suited for processing biomedical text.
PubMedBERT [36]. PubMedBERT is a domain-specific language model, which is pre-trained from scratch on a large corpus of biomedical text. Unlike previous approaches that rely on continual pre-training, PubMedBERT is trained solely on in-domain text, allowing it to better capture the nuances and complexity of biomedical language. The pre-training process is carried out using a collection of approximately $14$ million PubMed abstracts, which amounts to $3.2$ billion words. By focusing exclusively on biomedical literature, PubMedBERT can learn a vocabulary that is highly representative of the domain, leading to improved performance on downstream biomedical NLP tasks.
GatorTron [37]. GatorTron is a trained LLM specifically designed to extract and capture patient characteristics from EHRs and to answer medical queries. Studies [37] have demonstrated that GatorTron performs better in clinical tasks than previous models, highlighting its significant application potential.
Med-PaLM [38]. Med-PaLM enhances the application of medical LLMs through timely adjustments to PaLM, showing promising prospects in the MultiMedQA benchmark tests and highlighting areas needing improvement to meet healthcare standards.
MedAlpaca [39]. MedAlpaca boosts model performance significantly by fine-tuning the open-source AI model LLaMa with specialized training datasets. MedAlpaca aims to facilitate research and development in medical imaging, ultimately contributing to advancements in patient care and medical research.
LLaVA-Med [40]. Using cost-effective learning strategies, LLaVA-Med performs better in addressing biomedical imaging problems. This model demonstrates prowess in medical image understanding and analysis, capable of inferring patients’ pathological conditions from imaging data such as CT scans and X-rays. Equipped with multimodal dialogue capabilities, LLaVA-Med operates at the intersection of vision and language processing, enabling it to interpret visual medical data within a contextual framework.

### v2

@@ III-C Evaluations for Med-LLMs
III-C1 Quantitative Evaluations.
As shown in Table V, different tasks are evaluated via various metrics, covering accuracy, F1 score, and perplexity. Here, we introduce the details of some important metrics.
$\bullet$ Perplexity. Perplexity (PPL) is a standard that evaluates how well a probability model can predict a sample. When applied to language models like GPT, it represents the exponential average negative log-likelihood of a sequence. In essence, a lower perplexity score suggests that the model has a higher certainty in its predictions.
@@
$\bullet$ ROUGE. ROUGE scores compare the overlapping units (n-grams, skip-bigrams, or longest common subsequences) between the generated text and the reference. There are many ROUGE variants to measure the overlaps from different viewpoints, such as ROUGE-N, ROUGE-L, ROUGE-W, and ROUGE-SU. Here, we only give a detailed explanation for the most basic ROUGE-N. ROUGE-N is a metric that quantifies the overlapping N-grams between a candidate summary and a reference summary. Let $C$ be the set of all unigrams (or bigrams for ROUGE-2) in the candidate summary and $R$ be the set of all unigrams (or bigrams) in the reference summary. ROUGE-1 can be calculated as:
$P_{\text{ROUGE-1}}$ $=\frac{\sum_{w\in C}\text{Count}{R}(w)}{|C|},$ (2)
$R_{\text{ROUGE-1}}$ $=\frac{\sum_{w\in C}\text{Count}{R}(w)}{|R|},$
$\bullet$ BLEU. BLEU calculates the similarity between a candidate translation and one or more reference translations based on n-gram overlap. The idea is that a good translation should contain many n-grams (sequences of words) that are also present in the reference translations. Mathematically, the BLEU score is computed as follows:
@@ III-C2 Qualitative Evaluations
TABLE V: Commonly Used Evaluation Metrics for NLP tasks.
Task Evaluation Metrics
Text Generation BLEU, METEOR, ROUGE, CIDEr, Perplexity
Machine Translation BLEU, Translation Edit Rate, METEOR, Adequacy, Fluency
@@ III-C3 Evaluation Benchmark for LLMs’ Clinical Capabilities
$\bullet$ MedBench [17]. MedBench is a comprehensive evaluation benchmark for the medical domain, which consists of 40,041 questions about various branches of medicine including authentic examinations and clinical reports. The main components of MedBench involve: Chinese Medical Licensing Examination, Resident Standardization Training Examination, Doctor In-Charge Qualification Examination, and real-world clinic cases encompassing examinations, diagnoses, and treatments. Extensive experiments provided the following findings: (1) Chinese medical LLMs (e.g., HuaTuoGPT [159]) underperform on the MedBench, indicating the importance of clinical knowledge and diagnostic precision. (2) Some general LLMs (e.g., GPT 4) surprisingly convey considerable clinical knowledge.
@@
$\bullet$ AutoEval [29]. AutoEval is an automated evaluation framework to evaluate the realistic capabilities of LLMs, i.e., working as virtual doctors for multi-turn dialogue. The clinical demands for consultation tasks require LLMs to (1) consider the factors they do not know, (2) query about missing medical factors of patients, and (3) generate diagnosis and treatment results. To this end, AutoEval is to reformulate USMLE into medical multiple-choice questions. Then, several metrics are designed to indicate the ability of doctor LLMs in multi-turn consultation, i.e., Medical Information Coverage Rate, Accuracy of the Final Task, Average Turn, and Average Length.
Figure 4: Making a LLM to be a Doctor: A Multi-Step Approach. Prompt engineer can craft suitable prompts to derive the desired responses. Medical-specific fine-tuning can update parameters of a pre-trained LLM on a medical datasets to improve the clinical performance. RAG is to combining prompt engineering with context retrieval from external medical documents.
@@
$\bullet$ LLM-Mini-CEX [30]. Based on MiniCEX [174, 175], an evaluation criterion called LLM-specific Mini-CEX is established to effectively assess the clinical capabilities of LLMs. Then, a patient simulator is developed to simulate the conversations with LLMs, where ChatGPT is utilized to automatically evaluate diagnosis dialogues.
@@
$\bullet$ MedGPTEval [31]. MedGPTEval contains Chinese medical datasets and public benchmarks. The evaluation metrics are designed based on a comprehensive literature review, including medical professional capabilities, social comprehensive capabilities, contextual capabilities, and computational robustness, with 16 detailed indicators. These metrics have been optimized by a Delphi method using 5 experts in medicine and engineering [31]. Then, 3 clinical experts manually construct a set of medical datasets for interactions, including 7 case reports and 27 medical dialogues.
@@
$\bullet$ LLM-Human Evaluation [176]. LLMs have exhibited superior performance on unseen tasks, even only trained on the specific-task instructions. The study [176] investigates that if such an ability can be utilized as an alternative to human evaluation. Both human and LLM evaluations are conducted for two NLP tasks: adversarial attacks and open-ended story generation. The experiments indicate that LLM evaluation exhibits consistent results with human-level evaluation. Besides, LLM evaluation provides more stable results over different task instructions and the sampling algorithms for generation.
@@
$\bullet$ RJUA-SPs [32]. The RJUA-SPs evaluation approach contains three basic elements: (1) Metric. Professional clinical practice pathways are exploited to define the clinical capabilities that a doctor should possess, named LLM-specific clinical pathway (LCP). (2) Data. Standardized Patients (SPs) are utilized to guide the data collection, which aims to well maintain the completeness of the evaluation. (3) Algorithm. A multi-agent framework is formulated to simulate the interactive environment between SPs and a doctor agent. A retrieval-augmented algorithm is specially designed to measure whether the behaviors of a LLM doctor are consistent with LCP.
III-D Making a LLM to be a doctor
Transforming a general large language model (LLM) into a specialized medical LLM, akin to training it to become a “doctor”, involves several sophisticated steps that go beyond simple parameter adjustments. As shown in Figure 4, the overall process contains a multi-step approach: prompt engineering, medical-specific fine-tuning, and RAG.
$\bullet$ Prompt Engineering. This step involves crafting and refining prompts that can effectively guide the model to produce medically accurate and contextually appropriate outputs. The process begins with developing highly specific prompts, including detailed instructions, examples, and constraints that align closely with the medical domain. For instance, prompts might be designed to ask for diagnoses based on symptoms, treatment options for specific conditions, or explanations of complex medical procedures in layman’s terms. It’s crucial to incorporate elements such as patient history, current medications, and relevant clinical guidelines within the prompts to ensure the personalized responses.
$\bullet$ Medical-Specific Fine-tuning. The fine-tuning begins with selecting a high-quality dataset that encompasses a wide array of medical knowledge, including clinical notes, textbooks, research articles, patient records (with proper anonymization), and other relevant materials. The data should cover various aspects of medicine such as pathology, pharmacology, anatomy, diagnostics, treatment protocols, and patient care practices. Ensuring diversity in the dataset helps the model generalize better across different medical scenarios and specialties. Moreover, the fine-tuning process should include continuous learning mechanisms that allow the LLM to stay updated with the latest medical research and guidelines. This can be achieved through periodic retraining with new data or by implementing online learning techniques where the model can learn from new inputs without forgetting previously learned information—a challenge known as catastrophic forgetting.
$\bullet$ Medical-Specific RAG. Additionally, implementing a medical RAG requires integrating robust retrieval mechanisms that allow the model to access and leverage up-to-date, evidence-based medical information. RAG enhances the LLM’s capabilities by combining its generative power with precise, contextually relevant data retrieved from trusted medical databases and literature. This hybrid approach ensures that the outputs are not only coherent and contextually appropriate but also grounded in the latest clinical guidelines, research findings, and patient care standards. To achieve this, the system must be equipped with an efficient retrieval component capable of understanding complex medical queries and retrieving pertinent documents or passages from a vast corpus of medical knowledge. This corpus could include peer-reviewed articles, clinical trial results, textbooks, and practice guidelines. The retrieval process should prioritize high-quality sources and be able to handle nuanced queries related to diagnosis, treatment options, drug interactions, and patient-specific considerations.
Last but not least, rigorous testing through simulated clinical scenarios and peer review by medical professionals helps validate the model’s efficacy and reliability in real-world applications. Through these methods, a general LLM can be effectively adapted to serve as a powerful tool for supporting healthcare providers and improving patient outcomes.
III-E Specific Med-LLMs
@@
Existing LLMs in the medical field, such as Med-PALM, Codex-Med, and MedAlpaca, have contributed to advancements in healthcare with their unique design objectives, architectures, and functionalities. As shown in Figure 3, these models collectively drive the development of Med-LLMs and enhance their application in healthcare, providing robust support for future medical practice and research. Due to the limited space, we only simply introduce some Med-LLMs as follows.
  • BioBERT [143]: BioBERT is a variant of BERT specifically trained on biomedical text. It is designed to understand and process medical and biological information more effectively than general-purpose models. BioBERT is particularly useful for tasks such as named entity recognition, relation extraction, and text classification in the biomedical domain.
  • PubMedBERT [36]: PubMedBERT is a BERT-based model trained on PubMed abstracts. It is particularly effective for tasks involving biomedical literature, such as literature review, information retrieval, and summarization. PubMedBERT helps researchers quickly sift through large volumes of scientific papers to find relevant information.
  • ClinicalT5 [33]: ClinicalT5 is a T5-based model trained on clinical data. It is designed to handle clinical text generation and summarization tasks. ClinicalT5 can generate detailed clinical reports, summarize patient records, and assist in clinical decision-making.
  • GatorTron [81]: GatorTron is a large language model developed for various tasks, including those in the medical domain. It is known for its ability to handle complex medical text and generate high-quality outputs. GatorTron can be used for tasks such as generating detailed medical reports and summarizing clinical data.
  • Codex-Med [148]: Codex-Med is a variant of the Codex model, specifically tailored for medical applications, which can assist in generating detailed medical reports and summarizing clinical data.
  • Galactica [149]: Galactica is a large language model designed to handle a wide range of tasks, including those in the medical field. It is particularly effective for generating and understanding complex scientific and medical texts. Galactica can be used for literature review, information retrieval, and summarization in the biomedical domain.
  • Med-PaLM [38]: Med-PaLM is a language model specifically designed for medical applications. It is trained on a large corpus of medical data and is used for tasks such as medical text generation and information extraction. Med-PaLM can generate detailed medical reports and assist in clinical decision-making.
  • GPT-4-Med [72]: GPT-4-Med is a variant of the GPT-4 model, optimized for medical applications. It is designed to handle complex medical text and generate high-quality medical content. GPT-4-Med can be used for generating detailed medical reports, summarizing clinical data, and assisting in clinical decision-making.
  • ChatDoctor [151]: ChatDoctor is a conversational AI model designed to assist in medical consultations. It can answer medical questions and provide preliminary advice based on symptoms. ChatDoctor is useful for triaging patients and providing initial guidance before a formal medical consultation.
  • BenTsao [153]: BenTsao is a language model designed for medical applications. It is trained to handle tasks such as medical text generation and information retrieval. BenTsao can generate detailed medical reports and assist in clinical decision-making.
  • PMC-LLaMA [154]: PMC-LLaMA is a language model trained on PubMed Central data. It is designed to handle tasks related to biomedical literature, such as text generation and information retrieval. PMC-LLaMA helps researchers quickly sift through large volumes of scientific papers to find relevant information.
  • BianQue [156]: BianQue is a language model designed for medical applications. It is trained to handle tasks such as medical text generation and information retrieval. BianQue can generate detailed medical reports and assist in clinical decision-making.
  • Med-PaLM 2 [157]: Med-PaLM 2 is an updated version of Med-PaLM, designed to improve performance in medical applications. It is trained on a larger corpus of medical data and is used for tasks such as medical text generation and information extraction. Med-PaLM 2 can generate more accurate and detailed medical reports.
  • GatorTronGPT [158]: GatorTronGPT is a variant of GatorTron specifically designed for medical applications. It is used for tasks such as medical text generation and information retrieval. GatorTronGPT can generate detailed medical reports and assist in clinical decision-making.
  • HuatuoGPT [159]: HuatuoGPT is a language model designed for medical applications. It is trained to handle tasks such as medical text generation and information retrieval. HuatuoGPT can generate detailed medical reports and assist in clinical decision-making.
  • Med-Flamingo [163]: Med-Flamingo is a language model designed for medical applications. It is trained to handle tasks such as medical text generation and information retrieval. Med-Flamingo can generate detailed medical reports and assist in clinical decision-making.
  • ShenNong-TCM-LLM [9]: ShenNong-TCM-LLM specializes in traditional Chinese medicine (TCM). This model is trained on extensive TCM texts, including ancient classics, modern research papers, and clinical practice guidelines. ShenNong-TCM-LLM supports practitioners in diagnosing and treating patients using TCM principles, helps in the formulation of herbal prescriptions, and aids in educating both practitioners and patients about TCM practices.
  • MedicalGPT [7]: MedicalGPT is a general-purpose medical language model designed to support a wide range of healthcare-related tasks. It is trained on diverse datasets that include clinical notes, medical textbooks, and public health resources. MedicalGPT can generate detailed medical reports, assist in clinical decision-making, and provide information for patient consultations. Its versatility makes it suitable for various roles within the healthcare system.
  • QiZhenGPT [167]: QiZhenGPT focuses on traditional Chinese medicine diagnostics and treatment planning. It is trained on a comprehensive dataset of TCM case studies, diagnostic criteria, and therapeutic methods. QiZhenGPT assists TCM practitioners in making accurate diagnoses, recommending appropriate treatments, and educating patients about preventive health measures.
  • Med-ChatGLM [168] : Med-ChatGLM integrates conversational capabilities with deep medical knowledge. It is designed to engage in meaningful dialogues about health and wellness, provide personalized advice, and support users in managing chronic conditions. Med-ChatGLM can also assist healthcare providers in monitoring patient progress and adjusting treatment plans based on ongoing feedback.
  • HuatuoGPT-II [8]: HuatuoGPT-II is an advanced version of HuatuoGPT, optimized for improved performance in medical applications. It leverages a larger training dataset and refined algorithms to enhance its ability to generate detailed medical reports, summarize clinical data, and assist in clinical decision-making. HuatuoGPT-II’s improvements make it a valuable tool for healthcare professionals seeking precise and reliable AI support.
  • Taiyi-LLM [170]: Taiyi-LLM is a language model specialized in traditional Chinese medicine. It is trained on historical texts, contemporary research, and clinical practice data. Taiyi-LLM supports TCM practitioners in diagnosing and treating patients, formulating herbal prescriptions, and educating patients about TCM theories and practices. Its expertise in TCM makes it a valuable resource for integrating traditional and modern medical approaches.
  • Zhongjing [171]: Zhongjing is a language model focused on traditional Chinese medicine, named after the renowned physician Zhang Zhongjing. It is trained on a rich corpus of TCM literature and clinical data. Zhongjing aids practitioners in making accurate diagnoses, recommending effective treatments, and educating patients about TCM principles. It serves as a bridge between traditional wisdom and modern healthcare practices.
  • Med-Gemini [172]: Med-Gemini is a dual-language model capable of handling both English and Chinese medical texts. It is trained on a bilingual corpus of medical literature, clinical notes, and patient records. Med-Gemini facilitates cross-cultural medical communication, supports bilingual healthcare environments, and assists in translating medical documents accurately. Its dual-language capability enhances its utility in international healthcare settings.